For the following functions:

(a) f(x)=(x+2)²(x-1)

(b) f(x)= (x²)/(x²-9)

(c) f(x)= (x²-2x+1)/(x+1)


i. Find the critical numbers off (if any);

ii.Find the open interval(s) on which the function. increasing or decreasing

iii. Apply the first derivative test to identify all relative extrema;

iv. Use graphing utility to confirm your results.


1
Expert's answer
2021-12-05T16:55:15-0500

critical point of a function is point where it function is not differentiable or its derivative is 0


a)

"f'(x)=2(x+2)(x-1)+(x+2)^2=3x(x+2)=0"

"f'(x)>0" at "x\isin (-\infin,-2)" - "f(x)" is increasing

"f'(x)<0" at "x\isin (-2,0)" - "f(x)" is decreasing

"f'(x)>0" at "x\isin (0,\infin)" - "f(x)" is increasing

"f(x)" has relative maximum at x = -2

"f(x)" has relative minimum at x = 0


b)

"f'(x)=\frac{ 2x(x^2-9)-2x^3}{(x^2-9)^2}=-\frac{ 18x}{(x^2-9)^2}=0"

function is not differentiable at "x=\pm 3"

"f'(x)>0" at "x\isin (-\infin,-3)\cup (-3,0)" - "f(x)" is increasing

"f'(x)<0" at "x\isin (0,3)\cup (3,\infin)" - "f(x)" is decreasing

"f(x)" has relative maximum at x = 0


c)

"f'(x)=\frac{ x^2-2x+1}{x+1}=\frac{ (2x-2)(x+1)-(x-1)^2}{(x+1)^2}=\frac{ (x-1)(x+3)}{(x+1)^2}=0"

function is not differentiable at "x= -1"

"f'(x)>0" at "x\isin (-\infin,-3)" - "f(x)" is increasing

"f'(x)<0" at "x\isin (-3,1)" - "f(x)" is decreasing

"f'(x)>0" at "x\isin (1,\infin)" - "f(x)" is increasing

"f(x)" has relative maximum at x = -3

"f(x)" has relative minimum at x = 1